OpenStudy (anonymous):

What is the standard form equation of the line passing through (–2, 8) with a slope of 2?

OpenStudy (anonymous):

y = mx + c

OpenStudy (anonymous):

i thought it was +b

OpenStudy (anonymous):

Where m = slope c = Y intercept

OpenStudy (anonymous):

No Prob....u can also take it as b

OpenStudy (anonymous):

oh okay, so how would i start off the equation?

OpenStudy (anonymous):

Start from point-slope form: y-8 = 2(x+2) Simplify: y-8 = 2x+4 Subtract 2x from both sides: -2x + y - 8 = 4 Add 8 to both sides: -2x + y = 12 Multiply both sides by -1 to get a positive coefficient on x: 2x - y = -12

OpenStudy (anonymous):

thannk you!

OpenStudy (anonymous):

welcome